A device used in radiation therapy for cancer contains 0.17 g of cobalt-60 (59.933 819 u). The half-life of this isotope is 5.27 yr. Determine the activity (in Bq) of the radioactive material.
Answer #1

half life = ln(2)/k

k = ln(2)/half life = ln(2)/(5.27 x 365 x 24 x 3600) = 4.170693152 x 10-9 /s

now no. of molecules = (0.17/59.933819) x 6.023 x 1023 = 1.708401062 x 1021

activity = 4.170693152 x 10-9 x 1.708401062 x 1021 = 7.12521661 x 1012 Bq